OpenStudy (anonymous):

URGENT PLLZ HELP ASAP a broadway theatre has 500 seats, divided into orchestra, main, and blacony seating. orchestra seats sell for $60, main seats for $50, and balcony seats for $30. if all the seats are sold, the gross revenue to the theatre is $22,800. if all the main and balcony seats are sold, but only half the orchestra seats are sold, the gross revenue is $19,800. How many are there of each kind of seat?

OpenStudy (anonymous):

o+m+b=500 60o+50m+30b=22800 and 30o+50m+30b=19800 does that help?

OpenStudy (anonymous):

so o=100 m=240 b=160???

OpenStudy (anonymous):

looks good to me...good job

OpenStudy (anonymous):

thx :)

OpenStudy (anonymous):

I'm glad you did the matrix work lol -- i's a been up all night trying to get that

OpenStudy (anonymous):

once i get the equation i can get it lol thx :)